OpenStudy (anonymous):

simplify (1+i)^8 (write in form a+bi)

OpenStudy (anonymous):

easy

OpenStudy (anonymous):

express 1+i in polar form

OpenStudy (anonymous):

1+i = \[\sqrt{2}e ^{ \frac{i \pi}{4}}\]

OpenStudy (anonymous):

then take the power

OpenStudy (anonymous):

\[(1+I)^8 = [ \sqrt{2}e ^{\frac{i \pi}{4}}]^8\]

OpenStudy (anonymous):

= \[(\sqrt{2})^8 e ^{8\frac{ i \pi} {4}}\]

OpenStudy (anonymous):

= \[2^4 e ^{2 \pi i}\]

OpenStudy (anonymous):

but \[ e^{2 \pi i } =1 \]

OpenStudy (anonymous):

so answer = 2^4 = 16